a. In the agarose gel electrophoresis practical class, you are instructed to prepare a 750 ml of 1X TAE buffer from 25X TAE stock buffer. Subsequently, you are requested to prepare 35 ml of 0.75% (w/v) agarose gel using 1X TAE buffer that you prepared . Show your calculation and describe how you prepare the agarose gel
1
Expert's answer
2020-12-03T05:58:36-0500

Formula used:

C1V1=C2V2

Here C1= Initial concentration=25X

V1= Initial volume=?

C2= Final concentration=1X

V2= Final volume=750 ml

25XV1=1X750

V1=750/25=30ml

Thus 30ml of the stock solution is mixed with 720ml of distil water to make 750 ml of 1XTAE buffer.


agarose required (m)= percentage of agarose(w/v)X volume required

m= 0.75%X35

m= (0.75/100)X35

m=0.0075X35

m=0.26gm agarose is required to make 35ml of 0.75% 1XTAE buffer.


Preparation of agarose gel:

  1. Take 0.26gm of agarose and dissolve it in 35ml of 1XTAE buffer.
  2. Swirl the mixture and microwave it until the agarose is completely dissolved.
  3. Add ethidium bromide (0.5microgram/ml) and swirl to mix the content.
  4. Allow it to cool for some time.
  5. Setup the casting apparatus of the electrophoresis unit.
  6. Put comb and pour the gel in the gel casting tray.
  7. Allow it to cool at room temperature.
  8. Fill TAE buffer in the electrophoresis unit (cover gel also).
  9. Now carefully remove the comb.